A flywheel in the form of a uniformly thick disk of radius 1.53 m has a mass of 32.6 kg and spins counterclockwise at 113 rpm . Calculate the constant torque required to stop it in 2.25 min .
Answer #1

Moment of inertia, I = 0.5 x 32.6 x 1.53^2 = 38.16 kg m^2

Angular velocity, w = 113 x 2 pi/60 = 11.83 rad/sec

Acceleration, a = (0 - 11.83)/(2.25 x 60) = - 0.088 rad/s^2

Torque, T = 38.16 x 0.088 = - 3.34 Nm

Comment in case any doubt please rate my answer....